Operations Report- May 12th

Crew 212 Operations Report 12-MAY-2019

SOL: 7

Name of person filing report: Paolo Guardabasso

Non-nominal systems: None

Notes on non-nominal systems: Nothing to report

Generator (hours run last night): 11:15

On (last night): 8:15 PM

Off (this morning): 7:30 AM

Solar— SOC 61% at 7:45 PM [From Director, a
check will be run earlier in the afternoon from now on, to
avoid crossing the 80% limit again).


Diesel Reading – 90 %


Propane Reading – 46% [correcting the 45% of

Ethanol Free Gasoline – 4 gallons

Water (Axillary tank) – 0 gallons

Water (static tank) – 100 gallons

Axillary to Static tank transfer – no

Gallons transferred: No

Water in GreenHab – 140 gallons

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 143547.8

Toilet tank emptied: No

Deimos rover used: no (out of order)


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 92.8h

Beginning charge:

Ending charge: 79%

Currently charging: Yes

Opportunity rover used: No

Hours: 59.1h

Beginning charge:

Ending charge:

Currently charging: Yes

Curiosity rover used: Yes

Hours: 95.1h

Beginning charge: 100%

Ending charge: 78%

Currently charging: Yes

Notes on rovers: Curiosity was taken for the Engineering EVA that was performed to free the mouse (see general notes and comments)

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where? Nothing to report

CrewCar used and why, where? Off Campus

General notes and comments: During the night, a small mouse was captured with one of the two traps in the Hab. The morning, after a scheduled EVA was concluded, an Engineering EVA was organised to free the mouse close to Cowboy’s Corner. Crew Engineer Paolo and Crew Journalist Zoe travelled for about 20 minutes with Curi to the location and freed the mouse in a bush. On the way back there was a small problem with the radios, but it was solved as it was determined that a button was incidentally pushed while mounting the rover. Today was a resting day for most of the crew, and we had birthday cookies for our Spacesuit engineer Camilo.

Summary of internet: Nothing to report.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: A mouse was found on the upper deck around 1:45AM; it was inside one of the rooms and then moved to another, passing under the doors. A trap was placed and the mouse walked into it. An Engineering EVA was performed in the morning to release it. Other points: it was not possible to fix the oven; the Space chess board is missing a pawn; the leak from the pump has now penetrated the roof in the commander’s room through the loft floor plywood.

Summary of GreenHab operations: some sage was collected for lunch. A lizard was found inside, but it escaped once it saw a crew member.

Summary of ScienceDome operations: cleaning was performed by the crew.

Summary of RAM operations: Drone cleaning, repairs and testing were performed.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report.

Questions, concerns and requests to Mission Support: we need a static water tank refill. We have run out of kitchen paper.

Copyright © The Mars Society. All rights reserved. | Main Site